• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name gilman

    Log entry time 00:02:40 on June17,2000

    Entry number 43976

    keyword=shift summary swing june 16, 2000

    Arrive to find 107 uA on target and runs take 12 minutes. Yikes!


    time run events comments
    15:57 2134 0.7 M stop at 16:07 - IOC lost, ~15 minutes to reboot
    16:35 2135 1.3 M good run
    16:55 2136 1.3 M "
    17:10 2137 1.3 M "
    17:25 2138 1.3 M "
    17:40 2139 1.3 M "
    17:56 2140 1.3 M "
    18:14 2141 0.3 M thunderstorm passing, induces electrical noise
    causing RF trips, beam off and on a lot this run
    19:15 AES still investigating, 1 zone with low cooling water fault, not coming
    back up, +20 minutes from when it recovers...
    20:00 problem remains with low conductivity water system, expert on way in,
    magnetsd need to be cycled, expect at least 45 min -> 1 hour
    20:40 beam back, on and off...
    20:40 2142 test raster ...
    20:54 2143 1.3 M good run
    21:10 2144 1.2 M ", stop early since cavity lost at 21:25
    21:32 beam back...
    21:34 2145 1.3 M "
    21:54 2146 1.3 M "
    22:10 2147 1.3 M "
    22:28 2148 1.3 M "
    22:44 2149 0.9 M ", but roc14 crash at ~22:57?
    roc 14 gui reset + coda reset seem to work...
    2150 junk, coda bad - kill coda and restart on adaqs2 -
    it had been going on adaqs3
    23:11 2151 1.3 M "
    23:28 2152 1.3 M "
    23:45 2153 - still going at end of shift



    Current down to 95 microamps at end of shift.
    beam position on target not as stable as we would like, but within specs,
    called MCC about it several times..